OpenStudy (anonymous):

Simplify 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2. 2 square root of 8 8 square root of 2 8 square root of 6 6 square root of 8

OpenStudy (anonymous):

\[4\sqrt{2}+7\sqrt{2}-3\sqrt{2}\]

OpenStudy (anonymous):

The equation ^-^

OpenStudy (anonymous):

I am pretty sure it is 

TheSmartOne (thesmartone):

\[\sf 4\sqrt{2}+7\sqrt{2}-3\sqrt{2} = (4+7-3) \sqrt2\]

OpenStudy (anonymous):

So the answer is \[8\sqrt{2}\]

OpenStudy (anonymous):

Correct?

TheSmartOne (thesmartone):

correct (:

OpenStudy (anonymous):

Thank you!:)